# Spectral deviation of concentration operators on reproducing kernel Hilbert spaces

### 11 Mar 2026 v1

@@ 1.2. Time-frequency localization
where the integral converges in quadratic mean. The range-space of the short-time Fourier transform space $\mathbb{H}=V_{g}L^{2}(\mathbb{R}^{d})$ is a reproducing kernel subspace of $L^{2}(\mathbb{R}^{2d})$. For a domain $\Omega\subset\mathbb{R}^{2d}$, the concentration operator (1.1) implements the following time-frequency filter:
@@
While perfect time-frequency restriction to $\Omega$ is impossible due to the uncertainty principle, the time-frequency localization operator (1.4) was introduced in [13] as a suitable approximate restriction operator. The deviation of the spectral distribution function
(1.5) $\operatorname{trace}(1_{(\delta,1]}T_{\Omega})=\#\{\lambda\in\sigma(T_{\Omega}):\lambda>\delta\}$
@@
The concentration operator (1.4) is a Toeplitz operator with an indicator function as symbol, and its spectral asymptotics are classically studied under increasing large isotropic dilations $\Omega\mapsto R\Omega$ valid asymptotically for a fixed spectral threshold $\delta$. Precise two-term Szegő asymptotics for (1.4) with Schwartz window $g$ and a $\mathcal{C}^{2}$ domain $\Omega$ were derived in [44] and show that
@@
where $\log^{*}(x)=\max\{1,\log(x)\}$, $\mathcal{H}(\partial\Omega)$ is the perimeter of $\Omega$ and $C_{g,\partial\Omega}$ is a constant that depends explicitly on the curvature of $\Omega$ (in a measure theoretic sense) and the time-frequency decay of $g$. Applying (1.7) to the scaled domain $R\Omega$ on obtains estimates compatible with (1.6), but valid in the two parameter regime — see also [39] for a more technical version of (1.7), with an improved dependence on $\beta$ in the dilation regime.
@@ 1.3. Spectral stability under discretization
The motivating question for our work is whether Gabor multipliers satisfy spectral deviation estimates analogous to (1.7), that are also stable under refinements of the sampling lattice $\Lambda$, and faithful to the continuous limit. In other words, can we see the spectral deviation estimate (1.7) in practice? This question is subtle and continues to receive attention from practitioners [23]. While the Gabor multiplier (1.9) converges in operator norm to the time-frequency localization operator (1.4), this fact is far from enough to imply (two-parameter) spectral deviation estimates for Gabor multipliers.
@@
where, as before, $C_{g,\partial\Omega}$ is a constant that depends explicitly on the curvature of $\Omega$ and the time-frequency decay of $g$. Importantly, the right-hand side of (1.10) does not depend on the discretization parameter $N$. We also obtain similar estimates for windows $g$ in modulation spaces [4] that provide a discrete analogue to [39, Theorem 1.4], without discretization artifacts.
@@ 1.4. Further applications
The goal of treating discrete and continuous settings simultaneously and uniformly naturally leads us to work with (1.1) in the more general context of reproducing kernel Hilbert spaces. Beyond time-frequency analysis, our work implies the preservation of certain spectral properties under discretization of general frames (via frame multipliers) and concerns also vector-valued contexts (relevant in quantum harmonic analysis). While, at first glance, the quality of our estimates directly depends on the off-diagonal decay of the associated reproducing kernel, it is possible to combine our main result with certain decomposition methods and also treat slowly decaying kernels. As a proof of concept, we shall re-interpret the wave-packet expansion of [28] as a means to decompose the sinc kernel into kernels that fall into the scope of this work; see [31, 32] for a related technique.
@@ 2.1. Setup
is finite. Let $\mathbb{H}\subseteq L^{2}(X,\mathcal{H})$ be a reproducing kernel Hilbert space (RKHS) of vector-valued functions with reproducing kernel $K\colon X\times X\to\mathcal{S}_{2}(\mathcal{H})$, where $\mathcal{S}_{p}=\mathcal{S}_{p}(\mathcal{H})$ are the Schatten $p$ operators on $\mathcal{H}$, write $K_{y}(x)=K(x,y)$ — see Section 3.2 for details.
We shall study the spectrum of the concentration operator (1.1). For technical reasons it will be convenient to extend $T_{\Omega}$ by 0 on $L^{2}(X,\mathcal{H})\ominus\mathbb{H}$, that is,
@@ 2.2. Assumptions
    There exist a constant $\gamma>0$ — called inflation rate — such that
@@
$\bullet$ Condition [C2] controls the inflation rate $\gamma$ of the set $\Omega$ and its complement; the constant $\operatorname{\nabla}(\Omega)$ defined in (2.2) is called the inflation constant of $\Omega$. When $\Omega$ is subset of $\mathbb{R}^{d}$ with smooth boundary, then $\gamma=d$ and $\operatorname{\nabla}(\Omega)$ can be controlled in terms of the measure and curvature of $\partial\Omega$ (see Section 4). In other contexts, such as discrete or non-convex spaces, $\operatorname{\nabla}(\Omega)$ cannot be related to the topological boundary of $\Omega$. As we shall see, the inflation constant is compatible with natural approximation procedures, and that is its main merit for the problem under study.
$\bullet$ The doubling condition [C3] is easy to check in many examples and let us prove stronger estimates.
@@ 2.5.1. Applicability
(2.9) $\operatorname{\nabla}(\Omega)\lesssim\mathcal{H}(\partial\Omega)\,.$
@@ 2.6.1. Gabor multipliers
Let us state how Theorem 2.2 applies to Gabor multipliers (see Section 8 for definitions). To avoid technicalities, we restrict to dimension 1; in higher dimensions and for arbitrary full-rank lattices $\Lambda$, the eccentricity of $\Lambda$ and the regularity of $\partial\Omega$ come into play (see Theorem 8.3).
@@
The strength of these estimates lies in the fact that the right-hand sides do not depend on any discretization parameter (in particular, they do not involve the so-called dual Gabor window).
@@ 2.6.3. The Fourier transform and slow decaying kernels
The two-parameter spectral deviation estimates for time-frequency localization operators discussed in Section 1.2 parallel current developments in the study of concentration operators of the Fourier transform. Here, the relevant function space is the space of bandlimited functions, that is, square-integrable functions with Fourier transform supported on a given compact interval of the real line. One-parameter spectral asymptotics go back to Landau and Widom [34, 37] and are a building block of signal processing applications [48, 35, 36, 51, 54]. Motivated by modern applications, two-parameter estimates for Fourier concentration operators have been recently developed in [30, 45, 28, 7, 7, 14, 31, 32], while the case of higher dimensional general domains was treated in [29, 41, 26, 27].
@@ 2.6.4. Quantum harmonic analysis
Time-frequency localization has also been studied in vector-valued contexts [38], often under the name of quantum harmonic analysis. In Section 9.1 we show how Theorem 2.2 applied to so-called mixed-state localization operators greatly improving the results of [38, Theorem 4.4 and Lemma 5.3] in terms of the dependence on the spectral parameter.
@@ 3.1. Notation
We write $a\lesssim b$ whenever $a\leq Cb$ for some constant $C>0$ possibly depending on the parameters $\alpha,\beta,B$ (from (2.7), (8.2)), but not on any others. Similarly, we write $a\simeq b$ whenever $a\lesssim b\lesssim a$. Also, we will make use of the notation $\log^{\ast}(x):=\max\{1,\log x\}.$ When we need to stress the dependence of a quantity on a certain parameter, we add additional sub-indices. For example, we may write (2.4) as $N_{\mathbb{H}}(s)$.
@@ 3.2. Vector-valued reproducing kernel Hilbert spaces
In this section, we recall the definition and some basic properties of vector-valued reproducing kernel Hilbert spaces (RKHS). For a thorough introduction to that matter we refer to [46, Chapter 6]. Let $(X,\mu)$ be a measure space and $\mathcal{H}$ be a separable Hilbert space. We consider a linear space $\mathbb{H}\subset\mathcal{H}^{X}$ of functions from $X$ to $\mathcal{H}$ with the following properties:
@@ 4.2. Perimeter in Poincaré spaces
We start with a few definitions.
A curve is a continuous map $\sigma:[a,b]\to X$ with $a,b\in\mathbb{R}$. Its length is defined as
@@ 4.4. Non-local perimeters
We now study expressions of the form $\int_{\Omega}\int_{\Omega^{c}}\varphi(x,y)\,dx\,dy$, known as non-local perimeters, in terms of $\Omega$ and the off-diagonal decay of $\varphi$. Estimates of this kind are well-known for $\mathbb{R}^{d}$, and usually stated for convolution kernels $\varphi(x-y)$ (e.g. [42, Proposition 1.4]). We derive variants of such estimates for rather general spaces.
@@
In this case, by the doubling assumption [C3] for every $k\geq 0$,
@@ 5.1. Perimeter in discrete grids
(5.1) $\mu_{\Lambda}(\{\lambda\})=|\Lambda|:=\det A.$
@@
where $A(j)\in\mathbb{R}^{(d-1)\times d}$ is obtained from deleting the $j$-th column of $A$. By Cauchy-Binet,
@@ 5.2. Stability under discretization
(5.13) $\operatorname{\nabla}(\Omega_{\Lambda})$ $\lesssim\max\{1,\|A\|^{d}\}\frac{\Theta_{\omega}}{|\Lambda|}\frac{\mathcal{H}(\partial\Omega)}{\kappa\,\min\{1,\eta^{d-1}\}}\,,\qquad\text{and,}$
@@ 6.1. Schatten-norm estimates for Hankel operators
The next proposition provides a bound for the Schatten $p$-norms of the Hankel operator $H=H_{\Omega}=(1-P)1_{\Omega}P$. Crucially, we quantify the dependence on $p$ for $p\to 0^{+}$. With the set of tools developed in Section 4 at hand, we shall parallel and aptly adapt the argument of [39, Proposition 3.1]. Recall the dyadic decay measure $N(s)$ from (2.4).
@@
For $A=(H^{\ast}H)^{p/2}$, using [39, Lemma 2.1] we obtain
@@ 6.2. Proof of the main result
where $\tau=\max\{\tfrac{1}{\delta},\tfrac{1}{1-\delta}\}$; see, e.g., [39, Lemma 4.1] for details. Combining this with Proposition 6.1, we conclude that for $\alpha\in(0,1]$ and $p\in(0,2]$,
@@
(6.7) $\alpha=\log^{\ast}\Big((\tau{}N(s))^{\tfrac{1}{s}}\Big)^{-1}\,,\qquad\text{and}\qquad p=2\frac{\gamma+\alpha}{s+\alpha}\,$,
@@ 7. Toeplitz operators and frame multipliers
Applying our main results Theorem 2.2 and Corollary 2.3 will allow us to derive the following spectral deviation bounds for frame multipliers. A key quantity in our estimates is $\omega_{\lambda}:=\langle\varphi^{\mathrm{d}}_{\lambda},\varphi_{\lambda}\rangle$, the diagonal of the cross-Gram matrix of $\varphi$ and $\varphi^{\mathrm{d}}$. The mapping $\omega$ defines $\mu_{\varphi}:=\mu_{\omega}$ via (5.4), together with the extrema $\vartheta_{\varphi}$ and $\Theta_{\varphi}$. Moreover, we set $C_{\varphi}:={\Theta_{\varphi}^{8}}/({\vartheta_{\varphi}^{7}\,|\Lambda|})$ and note that
@@
If $\{\varphi_{\lambda}\}_{\lambda\in\Lambda}$ is a tight frame (that is, $a=b$ in the frame inequality (7.1)) consisting of unit norm elements, then Theorem 7.1 simplifies and yields a bound for the spectral deviation from $|\Omega|$ itself. Indeed, the dual frame is simply $\varphi^{\mathrm{d}}_{\lambda}=a^{-1}\varphi_{\lambda}$, which gives $\omega_{\lambda}\equiv a^{-1}$. Then,
@@ 8.1. Time-frequency concentration operators
Recall the time-frequency filter $A_{\Omega}^{g}$ from (1.4). Heuristically, $A_{\Omega}^{g}$ is approximately a projection onto the space of functions whose short-time Fourier transforms are mainly localized on $\Omega$. Theorem 2.2 and Corollary 2.3, which, essentially recover the main results of [39], help validate such intuition.
@@
The STFT is an isometric isomorphism $V_{g}:L^{2}(\mathbb{R}^{2d})\to\mathbb{H}_{g}$, and, with respect to the decomposition $L^{2}(\mathbb{R}^{2d})=\mathbb{H}_{g}\bigoplus\mathbb{H}_{g}^{\perp}$, the Toeplitz and localization operators are related by

### 08 Apr 2026 v2

@@ 1.2. Time-frequency localization
where the integral converges in quadratic mean. The range-space of the short-time Fourier transform $\mathbb{H}=V_{g}L^{2}(\mathbb{R}^{d})$ is a reproducing kernel subspace of $L^{2}(\mathbb{R}^{2d})$. For a domain $\Omega\subset\mathbb{R}^{2d}$, the concentration operator (1.1) implements the following time-frequency filter:
@@
While perfect time-frequency restriction to $\Omega$ is impossible due to the uncertainty principle, the time-frequency localization operator (1.4) was introduced in [13] as a suitable approximate restriction operator. Except for the multiplicity of the eigenvalue 0, the spectra of $T_{\Omega}$ and $A_{\Omega}^{g}$ coincide. The deviation of the spectral distribution function
(1.5) $\operatorname{trace}(1_{(\delta,1]}A_{\Omega}^{g})=\operatorname{trace}(1_{(\delta,1]}T_{\Omega})=\#\{\lambda\in\sigma(T_{\Omega}):\lambda>\delta\}$
@@
The concentration operator $T_{\Omega}$ is a Toeplitz operator with an indicator function as symbol, and its spectral asymptotics are classically studied under increasingly large isotropic dilations $\Omega\mapsto R\Omega$ valid asymptotically for a fixed spectral threshold $\delta$. Precise two-term Szegő asymptotics for (1.4) with Schwartz window $g$ and a $\mathcal{C}^{2}$ domain $\Omega$ were derived in [45] and show that
@@
where $\log^{*}(x)=\max\{1,\log(x)\}$, $\mathcal{H}(\partial\Omega)$ is the perimeter of $\Omega$ and $C_{g,\partial\Omega}$ is a constant that depends explicitly on the curvature of $\Omega$ (in a measure theoretic sense) and the time-frequency decay of $g$. Applying (1.7) to the scaled domain $R\Omega$ one obtains estimates compatible with (1.6), but valid in the two parameter regime — see also [40] for a more technical version of (1.7), with an improved dependence on $\beta$ in the dilation regime.
@@ 1.3. Spectral stability under discretization
The motivating question for our work is whether Gabor multipliers satisfy spectral deviation estimates analogous to (1.7), that are also stable under refinements of the sampling lattice $\Lambda$, and faithful to the continuous limit. In other words, can we see the spectral deviation estimate (1.7) in practice? This question is subtle and continues to receive attention from practitioners [23]. While the Gabor multiplier (1.9) converges in operator norm to the time-frequency localization operator (1.4), this fact is insufficient to imply (two-parameter) spectral deviation estimates for Gabor multipliers.
@@
where, as before, $C_{g,\partial\Omega}$ is a constant that depends explicitly on the curvature of $\Omega$ and the time-frequency decay of $g$. Importantly, the right-hand side of (1.10) does not depend on the discretization parameter $N$. We also obtain similar estimates for windows $g$ in modulation spaces [4] that provide a discrete analogue to [40, Theorem 1.4], without discretization artifacts.
@@ 1.4. Further applications
The goal of treating discrete and continuous settings simultaneously and uniformly naturally leads us to work with (1.1) in the more general context of reproducing kernel Hilbert spaces. Beyond time-frequency analysis, our work implies the preservation of certain spectral properties under discretization of general frames (via frame multipliers) and concerns also vector-valued contexts (relevant in quantum harmonic analysis). While, at first glance, the quality of our estimates directly depends on the off-diagonal decay of the associated reproducing kernel, it is possible to combine our main result with certain decomposition methods and also treat slowly decaying kernels. As a proof of concept, we shall reinterpret the wave-packet expansion of [28] as a means to decompose the sinc kernel into kernels that fall into the scope of this work; see [31, 32] for a related technique.
@@ 2.1. Setup
is finite. Let $\mathbb{H}\subseteq L^{2}(X,\mathcal{H})$ be a reproducing kernel Hilbert space (RKHS) of vector-valued functions with reproducing kernel $K\colon X\times X\to\mathcal{S}_{2}(\mathcal{H})$, where $\mathcal{S}_{p}=\mathcal{S}_{p}(\mathcal{H})$ are the Schatten $p$ operators on $\mathcal{H}$, and write $K_{y}(x)=K(x,y)$ — see Section 3.2 for details.
We shall study the spectrum of the concentration operator (1.1). For technical reasons, it will be convenient to extend $T_{\Omega}$ by 0 on $L^{2}(X,\mathcal{H})\ominus\mathbb{H}$, that is,
@@ 2.2. Assumptions
    There exists a constant $\gamma>0$ — called inflation rate — such that
@@
$\bullet$ Condition [C2] controls the inflation rate $\gamma$ of the set $\Omega$ and its complement; the constant $\operatorname{\nabla}(\Omega)$ defined in (2.2) is called the inflation constant of $\Omega$. When $\Omega$ is a subset of $\mathbb{R}^{d}$ with smooth boundary, then $\gamma=d$ and $\operatorname{\nabla}(\Omega)$ can be controlled in terms of the measure and curvature of $\partial\Omega$ (see Section 4). In other contexts, such as discrete or non-convex spaces, $\operatorname{\nabla}(\Omega)$ cannot be related to the topological boundary of $\Omega$. As we shall see, the inflation constant is compatible with natural approximation procedures, and that is its main merit for the problem under study.
$\bullet$ The doubling condition [C3] is easy to check in many examples and allows us prove stronger estimates.
@@ 2.5.1. Applicability
(2.9) $\operatorname{\nabla}(\Omega)\lesssim\mathcal{H}(\partial\Omega)\,$,
@@ 2.6.1. Gabor multipliers
Let us state how Theorem 2.2 applies to Gabor multipliers (see Section 8 for definitions).
@@
The strength of these estimates lies in the fact that the right-hand sides do not depend on any discretization parameter (in particular, they do not involve the so-called dual Gabor window). The frame assumption of the Corollary is not a restriction, as it holds automatically for all large $N$. A version of Corollary 2.4 holds in higher dimension and for time-frequency parameters discretized along a general full-rank lattice $\Lambda$; in this setting, the eccentricity of $\Lambda$ and the regularity of $\partial\Omega$ come into play (see Theorem 8.3).
@@ 2.6.3. The Fourier transform and slow decaying kernels
The two-parameter spectral deviation estimates for time-frequency localization operators discussed in Section 1.2 parallel current developments in the study of concentration operators of the Fourier transform. Here, the relevant function space is the space of bandlimited functions, that is, square-integrable functions with Fourier transform supported on a given compact interval of the real line. One-parameter spectral asymptotics go back to Landau and Widom [35, 38] and are a building block of signal processing applications [49, 36, 37, 52, 55]. Motivated by modern applications, two-parameter estimates for Fourier concentration operators have been recently developed in [30, 46, 28, 7, 14, 31, 32], while the case of higher dimensional general domains was treated in [29, 42, 26, 27, 33].
@@ 2.6.4. Quantum harmonic analysis
Time-frequency localization has also been studied in vector-valued contexts [39], often under the name of quantum harmonic analysis. In Section 9.1 we show how Theorem 2.2 applies to so-called mixed-state localization operators greatly improving the results of [39, Theorem 4.4 and Lemma 5.3] in terms of the dependence on the spectral parameter.
@@ 3.1. Notation
We write $a\lesssim b$ whenever $a\leq Cb$ for some constant $C>0$ possibly depending on the parameters $\alpha,\beta,B$ (from (2.7), (8.2)), but not on any others. Similarly, we write $a\simeq b$ whenever $a\lesssim b\lesssim a$. Also, we will make use of the notation $\log^{\ast}(x):=\max\{1,\log x\}.$ When we need to stress the dependence of a quantity on a certain parameter, we add additional subscripts. For example, we may write (2.4) as $N_{\mathbb{H}}(s)$.
@@ 3.2. Vector-valued reproducing kernel Hilbert spaces
In this section, we recall the definition and some basic properties of vector-valued reproducing kernel Hilbert spaces (RKHS). For a thorough introduction to that matter we refer to [47, Chapter 6]. Let $(X,\mu)$ be a measure space and $\mathcal{H}$ be a separable Hilbert space. We consider a linear space $\mathbb{H}\subset\mathcal{H}^{X}$ of functions from $X$ to $\mathcal{H}$ with the following properties:
@@ 4.2. Perimeter in Poincaré spaces
We start with a few definitions. A curve is a continuous map $\sigma:[a,b]\to X$ with $a,b\in\mathbb{R}$. Its length is defined as
@@ 4.4. Non-local perimeters
We now study expressions of the form $\int_{\Omega}\int_{\Omega^{c}}\varphi(x,y)\,dx\,dy$, known as non-local perimeters, in terms of $\Omega$ and the off-diagonal decay of $\varphi$. Estimates of this kind are well-known for $\mathbb{R}^{d}$, and usually stated for convolution kernels $\varphi(x-y)$ (e.g. [43, Proposition 1.4]). We derive variants of such estimates for rather general spaces.
@@
In this case, by the doubling assumption [C3], for every $k\geq 0$,
@@ 5.1. Perimeter in discrete grids
(5.1) $\mu_{\Lambda}(\{\lambda\})=|\Lambda|:=|\det A|.$
@@
where $A(j)\in\mathbb{R}^{(d-1)\times d}$ is obtained by deleting the $j$-th column of $A$. By Cauchy-Binet,
@@ 5.2. Stability under discretization
(5.13) $\operatorname{\nabla}(\Omega_{\Lambda})$ $\lesssim\max\{1,\|A\|^{d}\}\frac{\Theta_{\omega}}{|\Lambda|}\frac{\mathcal{H}(\partial\Omega)}{\kappa\,\min\{1,\eta^{d-1}\}}\,,\qquad\text{and}$
@@ 6.1. Schatten-norm estimates for Hankel operators
The next proposition provides a bound for the Schatten $p$-norms of the Hankel operator $H=H_{\Omega}=(1-P)1_{\Omega}P$. Crucially, we quantify the dependence on $p$ for $p\to 0^{+}$. With the set of tools developed in Section 4 at hand, we shall parallel and aptly adapt the argument of [40, Proposition 3.1]. Recall the dyadic decay measure $N(s)$ from (2.4).
@@
For $A=(H^{\ast}H)^{p/2}$, using [40, Lemma 2.1] we obtain
@@ 6.2. Proof of the main result
where $\tau=\max\{\tfrac{1}{\delta},\tfrac{1}{1-\delta}\}$; see, e.g., [40, Lemma 4.1] for details. Combining this with Proposition 6.1, we conclude that for $\alpha\in(0,1]$ and $p\in(0,2]$,
@@
(6.7) $\alpha=1/\log^{\ast}\big((\tau{}N(s))^{1/s}\big)\,,\qquad\text{and}\qquad p=2\frac{\gamma+\alpha}{s+\alpha}\,$,
@@ 7. Toeplitz operators and frame multipliers
Applying our main results Theorem 2.2 and Corollary 2.3 will allow us to derive the following spectral deviation bounds for frame multipliers. A key quantity in our estimates is $\omega_{\lambda}:=\langle\varphi^{\mathrm{d}}_{\lambda},\varphi_{\lambda}\rangle$, the diagonal of the cross-Gram matrix of $\varphi$ and $\varphi^{\mathrm{d}}$. Using $\omega$, we define $\mu_{\varphi}:=\mu_{\omega}$ via (5.4), together with the extrema $\vartheta_{\varphi}$ and $\Theta_{\varphi}$. Moreover, we set $C_{\varphi}:={\Theta_{\varphi}^{8}}/({\vartheta_{\varphi}^{7}\,|\Lambda|})$ and note that
@@
If $\{\varphi_{\lambda}\}_{\lambda\in\Lambda}$ is a tight frame (that is, $a=b$ in the frame inequality (7.1)) consisting of unit-norm elements, then Theorem 7.1 simplifies and yields a bound concerning the spectral deviation from a multiple of $|\Omega|$. Indeed, the dual frame is simply $\varphi^{\mathrm{d}}_{\lambda}=a^{-1}\varphi_{\lambda}$, which gives $\omega_{\lambda}\equiv a^{-1}$, and,
@@ 8.1. Time-frequency concentration operators
Recall the time-frequency filter $A_{\Omega}^{g}$ from (1.4). Heuristically, $A_{\Omega}^{g}$ is approximately a projection onto the space of functions whose short-time Fourier transforms are mainly localized on $\Omega$. Theorem 2.2 and Corollary 2.3, which essentially recover the main results of [40], help validate such intuition.
@@
The STFT is an isometric isomorphism $V_{g}:L^{2}(\mathbb{R}^{d})\to\mathbb{H}_{g}$, and, with respect to the decomposition $L^{2}(\mathbb{R}^{2d})=\mathbb{H}_{g}\bigoplus\mathbb{H}_{g}^{\perp}$, the Toeplitz and localization operators are related by